Suppose you received an unexpected inheritance of 34,200. You decide to invest by placing some of the money in stocks and some in bonds. To diversify, you decide that four times the amount in bonds should equal five times the amount invested in stocks. How much should be invested in stocks? How much should be invested in bonds?
Answer by lwsshak3(11628)

Suppose you received an unexpected inheritance of 34,200. You decide to invest by placing some of the money in stocks and some in bonds. To diversify, you decide that four times the amount in bonds should equal five times the amount invested in stocks. How much should be invested in stocks? How much should be invested in bonds?
***
let x=amt invested in stocks
34200-x=amt invested in bonds
4(34200-x)=5x
136800-4x=5x
9x=136800
x=15200
34200-x=19000
amt invested in stocks=15200
amt invested in bonds=19000